OpenStudy (saaarrrahhhh):

What is the constant of variation, k, of the line y = kx through (3,18) and (5,30)?

myininaya (myininaya):

you can find the slope by doing \[\frac{y_2-y_1}{x_2-x_1}=\frac{30-18}{5-3}\]

OpenStudy (saaarrrahhhh):

@myininaya is that what the constant of variation is? the slope?